OpenStudy (anonymous):

y=1/5x i have to graph the y intercept

OpenStudy (amistre64):

the tint, is found at x=0

OpenStudy (amistre64):

.... yint, not tint lol

OpenStudy (anonymous):

youre graph would b HYPERBOLA when y=1/5x 5y=1/x 5xy=1 where x is not equal to 0 y=1/5x

OpenStudy (amistre64):

im assuming this is: \(\frac{1}{5}x\)

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

when x=0; (1/5)x = 0 ; so your yint = (0,0)

OpenStudy (anonymous):

you can see the graph:-)